OpenStudy (anonymous):

The perimeter of a rectangle is fifty centimeters. The length of the rectangle is twenty centimeters. What is the width of the rectangle? thirty centimetres fifteen centimetres ten centimetres five centimeters

OpenStudy (anonymous):

Perimeter = length + length + width + width P = 2 * l + 2 * w 50 = 2 * 20 + 2 * w

OpenStudy (anonymous):

w=*

OpenStudy (anonymous):

8

OpenStudy (anonymous):

50 = 40 + 2 * w 50 - 40 = 40 - 40 + 2 * w 10 = 2 * w

OpenStudy (anonymous):

10cm?

OpenStudy (anonymous):

10 = 2 * w 10 / 2 = 2 * w / 2 w = 5

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem!